OpenStudy (anonymous):

What is the possible value for the missing term of the geometric sequence? 1250, ... , 50, ... a.250 b.1200 c.125 d.650

OpenStudy (anonymous):

Let x be the missing number Then 1250/ x = x /50 x^2 = 1250 x 50 = 62500 x =250

OpenStudy (cwrw238):

a geometric sequence can be written as a , ar , ar^2 here a = 1250 and ar^2 = 50 so can we can find r from this info

OpenStudy (anonymous):

How did you get x=250?

OpenStudy (anonymous):

x = square root of 62500 = 250

OpenStudy (cwrw238):

ar^/a = r^2 = 50/1250 = 0.04 r^2 = 0.04 r = 0.2 so second term = ar = 1250 * 0.2 = 250

OpenStudy (anonymous):

Ohh so you take the square root?

OpenStudy (anonymous):

yes